Can somebody please help me out? I need help with sequence and series. Below is a problem that I need to solve using arithmetic series.
1)An embroidery pattern calls for five stitches in the first row and for three more stitches in each successive row. The 25th row, which is the last row, has 77 stitches. Find the total number of stitches in the pattern.

2) Then create your own sequence or series for your class to solve.

Answer by Theo(13342)

1)An embroidery pattern calls for five stitches in the first row and for three more stitches in each successive row. The 25th row, which is the last row, has 77 stitches. Find the total number of stitches in the pattern.

An = A1 + (n-1)*d

Sn = n*(A1 + An)/2

A1 = 5
d = 3

A25 = 5 + 24*3 = 77

S25 = 25*(5+77)/2 = 25*(82)/2 = 25*41 = 1025